A body is pushed giving it an initial velocity of 5 m/s and comes to rest after travelling 4.9m .find meu 
 

Grade:11

2 Answers

Arun
4 years ago
Let,u be initial velocity and v be final velocity.
s be displacement and a be the acceleration.
K be the coefficient of friction.
From kinematics,
v² = u² + 2as
0 = 52 - 2a(4.9) (since, it's decelerating)
a= 50/9.8 = 5.102
a = K.g = 0.8 (g is acceleration due to gravity, value = 10m/s²)
K = 5.102/9.8 = 0.521
Vikas TU
4 years ago
Dear student 
u = 5.6 m/sec
v = 0
s=8m
g=9.8m/sec^2
v^2 - u^2 = 2as
a= -5.6 * 5.6/16
a = -1.96 m/sec2
now,
coeff. of friction = a/g
= 1.96/9.8
= 0.2
Good Luck